Calculate the concentration of all species in a 0.155 M solution of H2SO3.


[H3O+] =_____M


[H2SO3]=_____M


[HSO3-]= _____ M


[SO3^2-]=_____M


[OH-]=______M
Answer #1

[H3O+] =___0.31__M


[H2SO3]=_0.155____M


[HSO3-]= ___0__ M


[SO3^2-]=___0.155__M


[OH-]=__0.155____M
Answer #2

[h3o+] = 2*.155 = .31

h2so3] = 0

[hso3-] = 0

[so3^2]- = .155

[oh]- = (10^-14)/.31 = 3.22*10^-14